OpenStudy (anonymous):

3y + 20 = 3 +2y (1 point) –1 over 17 72 over 3 23 –17

OpenStudy (solomonzelman):

3y + 20 = 3 +2y -2y -2y y + 20 = 3 -20 -20 y=?

OpenStudy (solomonzelman):

can you do it from here?

OpenStudy (anonymous):

yes thank you

OpenStudy (solomonzelman):

Alright, You welcome!